OpenStudy (anonymous):

vertex form of y=-2x^2 + 6x + 1

OpenStudy (anonymous):

do you know vertex form?

OpenStudy (anonymous):

a(x-h)^2 + k

OpenStudy (anonymous):

alright so we want to find (h,k) or the vertex y = ax2 + bx + c h=-b/2a , k=f(h)

OpenStudy (anonymous):

so we will find h first, what are a and b?